OpenStudy (anonymous):

Evaluate the following integral. integrate from -1 to 1 (x)/(x^2+2x+1)dx This is what I got so far using partial fractions I got = A/(x+1) + B/(x+1)^2 x = A(x+1)^2+B(x+1) not sure if this is right but I am stuck here.

OpenStudy (anonymous):

just A(x+1) + B = x

OpenStudy (anonymous):

remember you're, in effect, 'finding the CD' \[\frac{ A }{x+1} + \frac{ B }{ (x+1)^2 }\] \[\frac{ A }{x+1} *\frac{ (x+1) }{ (x+1) }+ \frac{ B }{ (x+1)^2 } = \frac{ x}{ (x+1)^2 } \]

zepdrix (zepdrix):

Ah yes good catch Alg! :D Your initial setup looks good max, but when you multiplied through by the denominator, those terms should have simplified down a little bit more.

OpenStudy (anonymous):

Thanks for the help thats the nudge I needed and I was able to solve the problem.